OpenStudy (anonymous):

The ratio of the perimeters of two rectangles is 4 to 9. The perimeter of the larger rectangle is 54 inches. What is the perimeter of the smaller rectangle? A. 36 inches B. 121.5 inches C. 24 inches D. 13.5 inches

OpenStudy (anonymous):

c is the correct answer

OpenStudy (anonymous):

if you would like to know how i got that then let me know

OpenStudy (sasogeek):

@HQuinn0016 please explain how you got your answer.... dishing out just answers can actually get you suspended.

OpenStudy (anonymous):

ya ya whatever. im not the only one that does it. but if you insist haha. 4/9=x/54 54*4/9=x 6*4=x 24=x

OpenStudy (mismycatmin):

he not right is A.36 because you mutiplie 4 by 9 and that gives you 36

OpenStudy (anonymous):

yes he is i got it right

OpenStudy (anonymous):

she******* haha